Weierstrass Product for the Gamma Function

L^{A}TEXis good for typesetting math. L^{A}TEXhandles the numbering for the enumerate list environment automatically. L^{A}TEX manages cross references automatically. The following formula may be found in Whittaker & Watson [4].

∫_{0}^{∞} t^{x}e^{−t} \frac{dt}{t} = \frac{1}{x} ∏_{k=1}^{∞} \frac{(1 + \frac{1}{k})^{x}}{(1 + \frac{x}{k})}

References

[1] Lars V. Ahlfors, Complex Analysis, 3rd edition, McGraw Hill, 1979.

[2] Leslie Lamport, L^{A}TEX: a Document Preparation System, 2nd edition, Addison-Wesley, 1994.

[3] Frank Mittelbach et al., The L^{A}TEXCompanion, 2nd edition, Addison-Wesley, 2004.

[4] E. T. Whittaker & G. N. Watson, A Course of Modern Analysis, 4th edition, Cambridge University Press, 1969.